OpenStudy (anonymous):

PLEASE HELP! Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + 1/2 (2x – 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined? 30 ml 35 ml 40 ml 45 ml

OpenStudy (anonymous):

You need to combine like terms so you can solve for x. Doing this you have x +2x-10+x-5 = 145 4x-15=145 4x=160 x=40 (amount of white paint) so, 2x-10 = 70 (amount of blue paint) 70-40 = 30 ml more blue paint was used than white paint